OpenStudy (anonymous):

Chapter 5:Indices and Logarithms Double checking. simplify the following: @ganeshie8

OpenStudy (anonymous):

\[5^{2n}-5^{2n+1}-5^{2n-1}\]

OpenStudy (anonymous):

\[5^{2n}-5^{2n}\times5^1-5^{2n}-5^{-1}\]

OpenStudy (anonymous):

\[5^{2n}(1-1\times5-\frac{ 1 }{ 5 })\]

OpenStudy (anonymous):

\[5^{2n}(1-5-\frac{ 1 }{ 5 })\]

OpenStudy (anonymous):

\[5^{2n}(-\frac{ 21 }{ 5 })\]

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

is my working correct?

ganeshie8 (ganeshie8):

looks good !

OpenStudy (anonymous):

Thnx @ganeshie8